1. Find the amount accumulated FV in the given annuity account. HINT [See Quick Example 1 and Example 1.] (Assume end-of-period deposits and compounding at the same intervals as deposits. Round your answer to the nearest cent.)

$150 is deposited monthly for 18 years at 2% per year

FV = $

2.

Find the periodic payments PMT necessary to accumulate the given amount in an annuity account. HINT [See Quick Example 2.] (Assume end-of-period deposits and compounding at the same intervals as deposits. Round your answer to the nearest cent.)

$50,000 in a fund paying 5% per year, with monthly payments for 5 years

PMT = $
Answer #1

12 0.02 -1225 here PMT=150, 8= 0.02 m=12,7:18 1 * FV PMT (11+1 ) +1) = 50woo (ore) (17 m.) = 5000 ( 0:525) T-0.779 Li2018 50%